# Acquired TASTE: Multimodal Stance Detection with Textual and Structural Embeddings

### v2

@@ Abstract
Stance detection plays a pivotal role in enabling an extensive range of downstream applications, from discourse parsing to tracing the spread of fake news and the denial of scientific facts. While most stance classification models rely on textual representation of the utterance in question, prior work has demonstrated the importance of the conversational context in stance detection. In this work we introduce TASTE – a multimodal architecture for stance detection that harmoniously fuses Transformer-based content embedding with unsupervised structural embedding. Through the fine-tuning of a pretrained transformer and the amalgamation with social embedding via a Gated Residual Network (GRN) layer, our model adeptly captures the complex interplay between content and conversational structure in determining stance. TASTE achieves state-of-the-art results on common benchmarks, significantly outperforming an array of strong baselines. Comparative evaluations underscore the benefits of social grounding – emphasizing the criticality of concurrently harnessing both content and structure for enhanced stance detection.
@@ 1 Introduction
Focusing on conversational stance, we are interested in detecting stance in a multi-participant conversation, both on the utterance and speaker level. To this end, we propose TASTE – a multimodal architecture combining Textual And STructural Embeddings. Multi-participant conversations unfold in tree structures that can be converted to speaker graphs (see details in Section 3.2). TASTE uses LLMs to represent utterances; The max-cut graph optimization problem is used to derive contextual node (speaker) embeddings from the conversation graph. The embeddings of the two modalities are fused through a GRN layer. We considered several other text and node representations as well as different ways to combine them in a principled way.
@@
We find that the heavy lifting is achieved by the conversation structure. Teasing text and structure modalities apart, we find that, in most cases, using the structure alone outperforms text-based models. Combining the textual modality with the structural further adds, on average, 12% to the accuracy. This result should not be read as “structure is more important than text”. Rather, we maintain that the stance signal is encoded in a structure with a higher signal-to-noise ratio compared to text. This makes sense as texts are produced in a conversational context by subjective individuals. We further discuss this idea in Section 7 as it goes back to the sociolinguistics concepts of face Goffman (1955), conversational norms Grice (1975), and communication grounding Clark and Brennan (1991).
Figure 1: Illustration of a conversation tree, its corresponding speaker network, and the derived user embeddings using the max-cut SDP. Node colors correspond to speakers. Tree nodes represent utterances; Network nodes represent speakers. Edge width corresponds to the number of interactions. Users’ embeddings lie on an $n$-dimensional sphere and are rounded to discrete values by projection onto a random hyperplane (dashed line).
@@ 2 Related Work
Modalities indicate the different mediums that information is conveyed through text (in natural language), code (formal language), audio, visual, etc. Using more than a single modality achieves an improvement in the performance of many ML systems on multiple tasks. In multimodal learning, the representations of the different modalities are used explicitly in the training phase Ramachandram and Taylor (2017); Xu et al. (2023), rather than just defining a feature vector, e.g., Aldayel and Magdy (2019); Lynn et al. (2019).
@@ 3.1 Task Definition
Given a set of authors $A$ participating in a discussion, the set $U^{a}=\{u^{a}_{1},u^{a}_{2},...,u^{a}_{n}\}$ denotes the utterances produced by $a\in A$ through the discussion, where each utterance $u^{a}_{j}$ and user $a$ have a label $y_{j,a},y_{a}\in\{+,-\}$ (pro, con).
In this work, we address two distinct stance prediction tasks: utterance-level and author-level. At the utterance level, the classification task is straightforward – learning a classifier $\sigma(u^{(a)}_{j},\cdot)=\hat{y}_{a,j}$ that minimizes some loss function $\ell(y_{j,a},\hat{y}_{a,j})$. The dot in $\sigma$ represents additional input that the classifier may take besides the utterance. In our case, this extra input is the speakers’ network graph (see Figure 1). In other cases that can be meta-data like the speaker’s age, gender, etc.
At the author/speaker level, we assume stance labels are assigned to speakers. This assumption is valid if a speaker holds a pre-formed and stable stance throughout the discussion. This is inherently the case in debate-like discussions like those in our datasets. The task, in this case, is to train a classifier $\sigma(U^{(a)},\cdot)=\hat{y}_{a}$, in which all the utterances of a user are fed to the classifier, plus additional data (the graph in our case).
@@ 3.2 Model Architecture and Components
The model depicted in Figure 2 is the utterance-level classifier. It receives as input two modalities: textual and structural. The textual modality is the embedding of the utterance, and the structural is an embedding of the speaker (details below). The two embeddings are then fused using a GRN unit and fed into an MLP classifier. The MLP outputs a vector that quantifies the likelihood of each tag $t$ in the tagset $T$ (in our case $T=\{+,-\}$), and the binary prediction is the tag with maximum likelihood.
The author’s stance is computed by feeding all her utterances into the model (Figure 2), getting the prediction for each utterance separately, and taking the majority vote.
@@
The representation of an utterance $u$ is derived directly from the pretrained model. Specifically, $\vec{u}$, the content embedding for $u$, is obtained as the [CLS] token vector, which is extracted by feeding $u$ to pretrained Sentence-BERT Reimers and Gurevych (2019).
@@
We follow Pick et al. (2022) in the structure representation and user embedding. The discussion tree is first converted to a conversation graph (see Figure 1); each node represents a speaker, and an edge $e_{ab}$ between nodes $a$ and $b$ indicates a direct interaction between them. $w_{ab}$ indicates the weight of $e_{ab}$, determined by the number and type of interactions. The edge weights in the interaction graph are calculated as:
$w_{a,b}$ $=\alpha(replies(a,b)+replies(a,b))$ (1)
$+\beta(quotes(a,b)+quotes(a,b))$
@@ 4 Data
(a) 4Forum dataset basic statistics.
@@ 5.1 Baselines
The STEM algorithm Pick et al. (2022) is a stance classification algorithm that uses only the conversation structure. At its core, it is based on the SDP spelled in Eq. (2) and has shown superior results in author-level stance detection across datasets. It proceeds in three steps, the first is to compute the 2-core of the conversation graph; next compute the SDP embedding on the 2-core and derive author classifications from the vectors; finally, propagate the labels to the non-core part of the graph in a greedy manner. The derivation of author labels from SDP vector embedding is done via hyper-plane rounding Goemans and Williamson (1995). The geometric positioning of the user vectors on the $n$-dimensional sphere is illustrated in Figure 1.
A simplified version of STEM, which we call for simplicity SDP, skips the 2-core computation and applies SDP to the entire graph. The labels are derived using the hyperplane rounding technique.
@@ 6 Results and Analysis
Using uni-modal approach, we can tease the content and the structure apart and gain some interesting insights about interplay between them. intense (dense) though short (in utterance length) interactions, such as those in the Abortion, Evolution, and Gay Marriage topics, are better captured by structural models such as the SDP, while less intense interactions that exhibit longer utterances, such as in Gun Control, are better captured by text based embeddings such as S-BERT. The uni-modal results along with the number of interactions and the number of tokens in each topic are provided in Table 3 in Appendix A.
@@
Integrating content and structural information in a multimodal manner proves robust across datasets, topics, and conversational dynamics. The multimodal approach grounds the potentially rich textual content in the relevant social context. An additional analysis of a concrete example is provided in Appendix B, underscoring the benefits of our multimodal approach.
@@
We considered a naive yet commonly used fusion strategy, namely concatenation of the representations (see results in Table 4 in Appendix C). Using a GRN layer significantly outperforms concatenation across all topics. This notable difference in the results can be attributed to the GRN layer’s ability to scrupulously combine the distinct properties of both content and structure embeddings. Unlike simple concatenation, which merely amalgamates or adds up the information, the GRN layer applies a more nuanced approach. It effectively ‘gates’ the information from each embedding, allowing for a selective integration process. This gating mechanism is particularly advantageous in handling smaller and noisier datasets where discerning relevant information is critical. The GRN layer’s capacity to attend to the the most pertinent features from both content and structural data results in a more accurate and robust stance detection model.
@@ 7 Discussion
Some recent work (see Section 2) use other, more recent, datasets. The focus of this work is the interplay between content and speaker dynamics, hence whese datasets cannot be used for proper comparison of the utterance level, nor for speaker-level stance.
@@
The work of Li et al. (2018), explicitly modeling text and structure, is the most similar to our work. However, while their strategy for integrating text, speaker, and structure can be viewed as early fusion [^2] We note that Li et al. do not refer to their model as multimodal and do not explicitly refer to the fusion strategy., resulting in global embeddings in one shared space, our methodology combines embeddings of the different modalities using joint (hierarchical) fusion strategy Huang et al. (2020); Xu et al. (2023). We believe that the joint fusion of separate embeddings provides flexibility that results in higher efficiency. Specifically, it allows the use of the SDP for speaker representation. These non-orthodox embeddings are learned efficiently at the conversation level in a completely unsupervised manner.
@@ 9 Limitations
Appendix A Uni-modal Results
Table 3 presents uni-modal results, along with the number of interactions and the number of tokens in each topic. The table promotes an interesting observation regarding the interplay between content and structure: intense (dense) though short (in utterance length) interactions, such as those in the Abortion, Evolution, and Gay Marriage topics, are better captured by structural models such as the SDP, while less intense interactions that exhibit longer utterances, such as in Gun Control, are better captured by text based embeddings such as S-BERT.
Abortion Evolution Gay Marriage Gun Control
# Interaction 14,460 9,640 13,764 6,462
Avg. tokens 115.5 157.8 112.7 140.3
S-BERT (Acc.) 67.7 69.8 68.9 71.7
SDP (Acc.) 81.2 76.3 78.1 67.7
Appendix B Analysis of an Illustrative Example
The exchange presented in Figure 3 provides an example for the value of the multimodal framework and the contribution of the SDP to understanding the global structure of a discussion. In this discussion on gun control, participants engage in a more complex interaction pattern than simple turn-taking, where each comment directly opposes the previous one. Instead, participants often respond supportively to others’ comments, creating a nuanced dynamic which locally violates the max-cut assumption, a violation that may also affect the tone of the comment. The first comment of C, which locally violates the max-cut hypothesis, is such an example of supportive utterance which reads quite ambiguous when taken out of context. However, by adding to the model the global overview of the SDP, our model can accurately capture these interactions and correctly classify the participants’ stances, circumventing the local max-cut violation and the diluted tone of the utterance. Diving into specific results, relying solely on SBERT led to a 71% accuracy for the entire conversation from which this excerpt is taken, while our best TASTE model achieved a 92% accuracy. This significant improvement underscores the effectiveness of combining textual and structural embeddings for nuanced stance detection.
Appendix C GRN and Naive Fusion

### v3

@@ Abstract
Stance detection plays a pivotal role in enabling an extensive range of downstream applications, from discourse parsing to tracing the spread of fake news and the denial of scientific facts. While most stance classification models rely on the textual representation of the utterance in question, prior work has demonstrated the importance of the conversational context in stance detection. In this work, we introduce TASTE – a multimodal architecture for stance detection that harmoniously fuses Transformer-based content embedding with unsupervised structural embedding. Through the fine-tuning of a pre-trained transformer and the amalgamation with social embedding via a Gated Residual Network (GRN) layer, our model adeptly captures the complex interplay between content and conversational structure in determining stance. TASTE achieves state-of-the-art results on common benchmarks, significantly outperforming an array of strong baselines. Comparative evaluations underscore the benefits of social grounding – emphasizing the criticality of concurrently harnessing both content and structure for enhanced stance detection.
@@ 1 Introduction
Focusing on conversational stance, we are interested in detecting stance in a multi-participant conversation, both on the utterance and speaker level. To this end, we propose TASTE – a multimodal architecture combining Textual And STructural Embeddings. Multi-participant conversations unfold in tree structures, which can be converted into speakers’ interactions graphs (see details in Section 3.2). TASTE uses LLMs to represent utterances; the max-cut graph optimization problem is used to derive contextual node (speaker) embeddings from the interaction graph. The embeddings of the two modalities are fused through a GRN layer. We considered several other text and node representations as well as different ways to combine them in a principled way.
@@
We find that the heavy lifting is achieved by the conversation structure. Teasing text and structure modalities apart, we find that, in most cases, using the structure alone outperforms text-based models. Combining the textual modality with the structural further adds, on average, 12% to the accuracy. This result should not be read as “structure is more important than text”. Rather, we maintain that the stance signal is encoded in a structure with a higher signal-to-noise ratio compared to text. This makes sense as texts are produced within a conversational context by subjective individuals. We further discuss this idea in Section 7 as it goes back to the sociolinguistics concepts of face Goffman (1955), conversational norms Grice (1975), and communication grounding Clark and Brennan (1991).
Figure 1: Illustration of a discussion tree, its corresponding speakers interactions graph, and the derived speakers embedding using the max-cut SDP. Node colors correspond to speakers. Tree nodes represent utterances; Graph nodes represent speakers. Edge width corresponds to the number of interactions. Speakers’ embeddings lie on an $n$-dimensional sphere and are rounded to discrete values by projection onto a random hyperplane (dashed line).
@@ 2 Related Work
Modalities indicate the different mediums that information is conveyed through: text (in natural language), code (formal language), audio, visual, etc. Using more than a single modality achieves an improvement in the performance of many ML systems on multiple tasks. In multimodal learning, the representations of the different modalities are used explicitly in the training phase Ramachandram and Taylor (2017); Xu et al. (2023), rather than just defining a feature vector, e.g., Aldayel and Magdy (2019); Lynn et al. (2019).
@@ 3.1 Task Definition
Given a set of authors $A$ participating in a discussion, the set $U^{a}=\{u^{a}_{1},u^{a}_{2},...,u^{a}_{n}\}$ denotes the utterances produced by $a\in A$ throughout the discussion, where each utterance $u^{a}_{j}$ and user $a$ have a label $y_{j,a},y_{a}\in\{+,-\}$ (pro, con).
In this work, we address two distinct stance prediction tasks: utterance-level and author-level. At the utterance level, the classification task is straightforward – learning a classifier $\sigma(u^{a}_{j},\cdot)=\hat{y}_{a,j}$ that minimizes some loss function $\ell(y_{j,a},\hat{y}_{a,j})$. The dot in $\sigma$ represents additional input that the classifier may take besides the utterance. In our case, this extra input is the speakers’ interactions graph (see Figure 1). In other cases that can be meta-data like the speaker’s age, gender, etc.
At the author/speaker level, we assume stance labels are assigned to speakers. This assumption is valid if a speaker holds a pre-formed and stable stance throughout the discussion. This is inherently the case in debate-like discussions like those in our datasets. The task, in this case, is to train a classifier $\sigma(U^{a},\cdot)=\hat{y}_{a}$, in which all the utterances of a user are fed to the classifier, plus additional data (the graph in our case).
@@ 3.2 Model Architecture and Components
The model depicted in Figure 2 is the utterance-level classifier. It receives as input two modalities: textual and structural. The textual modality is the embedding of the utterance, and the structural is an embedding of the speaker (details below). The two embeddings are then fused using a GRN unit, following a joint fusion strategy that dynamically combines the modalities during learning to better capture interdependencies between content and conversational structure. The fused representation is fed into an MLP classifier, which outputs a vector quantifying the likelihood of each tag $t$ in the tagset $T$ (in our case $T=\{+,-\}$), and the binary prediction is the tag with maximum likelihood.
@@
The author’s stance is computed by feeding all their utterances into the model (Figure 2), getting the prediction for each utterance separately, and taking the majority vote.
@@
The representation of an utterance $u$ is derived directly from the pretrained model. Specifically, $\vec{u}$, the content embedding for $u$, is obtained as the [CLS] token vector, which is extracted by feeding $u$ to pretrained Sentence-BERT (Reimers and Gurevych, 2019).
@@
We follow Pick et al. (2022) in the structure representation and user embedding. The discussion tree is first converted to an interactions graph (see Figure 1); each node represents a speaker, and an edge $e_{ab}$ between nodes $a$ and $b$ indicates a direct interaction between them. $w_{ab}$ indicates the weight of $e_{ab}$, determined by the number and type of interactions. The edge weights in the interaction graph are calculated as:
$w_{a,b}$ $=\alpha(replies(a,b)+replies(b,a))$ (1)
$+\beta(quotes(a,b)+quotes(b,a))$
@@ 4 Data
(a) 4Forums dataset basic statistics.
@@ 5.1 Baselines
The STEM algorithm Pick et al. (2022) is a stance classification algorithm that uses only the conversation structure. At its core, it is based on the SDP spelled in Eq. (2) and has shown superior results in author-level stance detection across datasets. It proceeds in three steps, the first is to compute the 2-core of the conversation graph; next compute the SDP embedding on the 2-core and derive author classifications from the vectors; finally, propagate the labels to the non-core part of the graph in a greedy manner. The derivation of author labels from SDP vector embedding is done via hyper-plane rounding Goemans and Williamson (1995). The geometric positioning of the speaker vectors on the $n$-dimensional sphere is illustrated in Figure 1 (right).
A simplified version of STEM, which we call SDP, for simplicity, skips the 2-core computation and applies SDP to the entire graph. The labels are derived using the hyperplane rounding technique.
@@ 6 Results and Analysis
Using uni-modal approach, we can tease the content and the structure apart and gain some interesting insights about interplay between them. Intense (dense) though short (in utterance length) interactions, such as those in the Abortion, Evolution, and Gay Marriage topics, are better captured by structural models such as the SDP, while less intense interactions that exhibit longer utterances, such as in Gun Control, are better captured by text based embeddings such as S-BERT. The uni-modal results along with the number of interactions and the number of tokens in each topic are provided in Table 3.
Abortion Evolution Gay Marriage Gun Control
# Interaction 14,460 9,640 13,764 6,462
Avg. tokens 115.5 157.8 112.7 140.3
S-BERT (Acc.) 67.7 69.8 68.9 71.7
SDP (Acc.) 81.2 76.3 78.1 67.7
@@
Integrating content and structural information in a multimodal manner proves robust across datasets, topics, and conversational dynamics. The multimodal approach grounds the potentially rich textual content in the relevant social context.
The exchange presented in Figure 3 provides an example for the value of the multimodal framework and the contribution of the SDP to understanding the global structure of a discussion. In this discussion on gun control, participants engage in a more complex interaction pattern than simple turn-taking, where each comment directly opposes the previous one. Instead, participants often respond supportively to others’ comments, creating a nuanced dynamic which locally violates the max-cut assumption, a violation that may also affect the tone of the comment. The first comment of C, which locally violates the max-cut hypothesis, is such an example of supportive utterance which reads quite ambiguous when taken out of context. However, by adding the global overview of the SDP to the model, it can accurately capture these interactions and correctly classify the participants’ stances, circumventing the local max-cut violation and the diluted tone of the utterance. Diving into specific results, relying solely on SBERT led to a 71% accuracy for the entire conversation from which this excerpt is taken, while our best TASTE model achieved a 92% accuracy. This significant improvement underscores the effectiveness of combining textual and structural embeddings for nuanced stance detection.
@@
We considered a naive yet commonly used fusion strategy, namely concatenation of the representations (see results in Table 4 in Appendix A). Using a GRN layer significantly outperforms concatenation across all topics. This notable difference in the results can be attributed to the GRN layer’s ability to scrupulously combine the distinct properties of both content and structure embeddings. Unlike simple concatenation, which merely amalgamates or adds up the information, the GRN layer applies a more nuanced approach. It effectively ‘gates’ the information from each embedding, allowing for a selective integration process. This gating mechanism is particularly advantageous in handling smaller and noisier datasets where discerning relevant information is critical. The GRN layer’s capacity to attend to the the most pertinent features from both content and structural data results in a more accurate and robust stance detection model.
@@ 7 Discussion
Some recent work (see Section 2) use other, more recent, datasets. The focus of this work is the interplay between content and speaker dynamics, hence these datasets cannot be used for proper comparison of the utterance level, nor for speaker-level stance.
@@
The work of Li et al. (2018), explicitly modeling text and structure, is the most similar to our work. However, while their strategy for integrating text, speaker, and structure can be viewed as early fusion [^2] We note that Li et al. do not refer to their model as multimodal and do not explicitly refer to the fusion strategy., where modalities are combined at the input level to produce global embeddings in one shared space, our methodology combines embeddings of the different modalities using joint (hierarchical) fusion strategy Huang et al. (2020); Xu et al. (2023), where interactions between modalities are dynamically modeled during the learning process. We believe that the joint fusion of separate embeddings provides flexibility that results in higher efficiency. Specifically, it allows the use of the SDP for speaker representation. These non-orthodox embeddings are learned efficiently at the conversation level in a completely unsupervised manner.
@@ 9 Limitations
Appendix A GRN and Naive Fusion